You kept the switch in position 2 for a long time.
Answer #1

(a) When a battery is connected to a series resistor and capacitor, the initial current is high as the battery transports charge from one plate of the capacitor to the other. The charging current asymptotically approaches zero as the capacitor becomes charged up to the battery voltage. Charging the capacitor storesenergy in the electric field between the capacitor plates. The rate of charging is typically described in terms of a time constant RC.

CV Charge on capacitor R Charging current 0 RC 2RC 3RC 4RC time

It is given that the battery voltage (V) = 9 V. Therefore (1-1/e) of battery voltage is: 9 x (1-1/e) = 5.69 V

Time taken for the voltage to increase to this value is 11.22 ms from the table.

Time constant = 11.22 ms.

(b) Time constant can be expressed in terms of resistance R and capacitance C as:

? = RC

or, 11.22 x 10-3 = 102 x C

or, C = 1.1 x 10-4 F = 110 ?F